Back to main Archive News page.
Friday 30 September 2005 - All systems ran without major problems
Thursday 29 September 2005 - All systems ran without major problems
Wednesday 28 September 2005 - DHCP problem and VOIP problem
09:00 A problem with the DHCP server is causing problems logging on this morning. ISS is working on the problem. Normal service will be resumed as soon as possible. 10:10 Update: The problem with the DHCP server has been fixed. If you have been having problems: 1) log on, 2) if that doesn't work - reboot and log on again, 3) if the second log on fails, call the ISS Service Desk x10987.
10:48 We are currently seeing some problems with the Voice over IP telephones (currently used in LEC, Infolab, IAS and ISS). Calls can be received but not made. We are investigating the problems and hope to have service restored shortly. 12:38 Update: The problem appears to be a hardware issue with the VOIP system's connection to the public telephone network. A fault has been raised with our support provider and we are awaiting replacement hardware.21:40 Update: Replacement hardware from the supplier was installed but proved to be faulty. We are now running on the original hardware that appears to be stable. The maintainance company are working to resolve the problems with the replacement.
Tuesday 27 September 2005 - Network configuration changes
07:00-09:00 Network redesign work took place in the Engineering, Biological Science, Environmental Science and LEC buildings, causing disruption to network services.
08:45-10:30 The Campus wireless network was subject to connectivity problems, which have now been fixed.
Work begins on Tuesday 27th Sept for the rewiring of Pendle computer lab. Work should be completed by Thursday afternoon 29th Sept.
Monday 26 September 2005 - Library PCs
Several of the older Library open-access PCs are being replaced over the next few days with new PCs with flat screen monitors.
Friday 23 September 2005 - Central Filestore (H: drive) maintenance, and ResNet problems
The problems on ResNet were fixed by 12:30.
The central filestore (H: drive) was unavailable from 23:00 to midnight, while a suspected software issue was investigated.
Thursday 22 September 2005 - ResNet Network problems
There are still ongoing problems on ResNet.
Wednesday 21 September 2005 - Network configuration changes & network problems
07:00-09:00 During the Library network at risk period, network redesign work took place in the Library building causing some disruption.
14:00-16:20 ISS performed a hardware upgrade on Departmental Filestore held
on CENTRAL-FILES (some departments map this as "G: Drive").
During this time Departmental Filestore was be available to read from, but not write to. Personal Filestore was unaffected by this change
Network problems: There were some network oddities in some areas of campus. The areas affected most appear to be Grizedale Lab however it is likely other parts of campus are without networking. The networking team are working to resolve this problem and service should be restored early tomorrow morning. Update 18:44: The bulk of the Residences Network (ResNet) is also affected. Service should be restored during the morning.
Tuesday 20 September 2005 - Network configuration changes
07:00-09:00 During the network at risk period, further work was carried out on the resdesign of the network, causing some disruption accross campus.
Monday 19 September 2005 - All systems ran without major problems
Friday 16 September 2005 - All systems ran without major problems
Thursday 15 September 2005 - All systems ran without major problems
Wednesday 14 September 2005 - CENTRAL-FILES upgrade
14:00-17:30 ISS performed a hardware upgrade on Personal Filestore held on CENTRAL-FILES ("H: Drive"). During this time Personal Filestore was available to read from, but not write to. Departmental Filestore was unaffected by this change.
Tuesday 13 September 2005 - Network configuration changes
07:00-10:00 Configuration changes were made to network equipment as part of the summer network infrastructure upgrades. This caused short network outages during the Tuesday 'At Risk' period to parts of campus.
Monday 12 September 2005 - All systems ran without major problems
Friday 09 September 2005 - All systems ran without major problems
Thursday 08 September 2005 - All systems ran without major problems
Wednesday 07 September 2005 - All systems ran without major problems
14:00-15:00 There were some very brief interruptions to the Exchange service to allow reconfiguration to match the new networking requirements.
Tuesday 06 September 2005 - All systems ran without major problems
07:30-08:30 The VPN Service was unavailable for a short time between for configuration of network settings.
Monday 05 September 2005 - All systems ran without major problems
During the week 29 August - 2 September, the PCs in Engineering A37 PC lab were upgraded. Unfortunately, the lab is closed until 30th September whilst major building work is undertaken.